# Selective Preference Optimization via Token-Level Reward Function Estimation

### v1

@@ Abstract
Recent advancements in large language model alignment leverage token-level supervisions to perform fine-grained preference optimization. However, existing token-level alignment methods either optimize on all available tokens, which can be noisy and inefficient, or perform selective training with complex and expensive key token selection strategies. In this work, we propose Selective Preference Optimization (SePO), a novel selective alignment strategy that centers on efficient key token selection. SePO proposes the first token selection method based on Direct Preference Optimization (DPO), which trains an oracle model to estimate a token-level reward function on the target data. This method applies to any existing alignment datasets with response-level annotations and enables cost-efficient token selection with small-scale oracle models and training data. The estimated reward function is then utilized to score all tokens within the target dataset, where only the key tokens are selected to supervise the target policy model with a reference model-free contrastive objective function. Extensive experiments on three public evaluation benchmarks show that SePO significantly outperforms competitive baseline methods by only optimizing 30% key tokens on the target dataset. SePO applications on weak-to-strong generalization show that weak oracle models effectively supervise strong policy models with up to 16.8$\times$ more parameters. SePO also effectively selects key tokens from out-of-distribution data to enhance strong policy models and alleviate the over-optimization problem.
@@ 1 Introduction
The recent development of large language models (LLMs) has focused on aligning model outputs with human intentions and preferences (Ji et al., 2023). During alignment, LLMs are directly optimized on pairwise data and response-level supervision, where popular methods such as reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022; Stiennon et al., 2020) and direct preference alignment Rafailov et al. (2024c); Yuan et al. (2023); Meng et al. (2024) only introduce supervision signals at the end of each response. As deriving preference optimization as bandit problems can lead to sub-optimal solutions and unstable training processes Zhong et al. (2024); Zeng et al. (2024), many works propose to model LLM decoding as token-level Markov Decision Processes (MDP) and introduce step-wise supervision signals that quantify the value of each action, successfully applied in tasks such as instruction following Zhong et al. (2024); Yoon et al. (2024) and mathematical reasoning Xie et al. (2024); Chen et al. (2024b); Lai et al. (2024).
@@
Figure 1: Token-level reward accumulations on two datasets. As tokens with high rewards are considered key tokens for chosen responses, their Top-Ktokens are accumulated in descending order with the highest rewards. In contrast, the rewards are accumulated in ascending orders for rejected responses.
@@
Though achieving outstanding performance, most of these methods are optimized on all available tokens from the training dataset. To validate the effectiveness of this setup, in Figure 1, we present the token-level reward accumulations for 1,000 samples from an instruction following dataset Cui et al. (2023) and a question answering (QA) dataset Wu et al. (2024), where the token-level rewards are assigned by GPT-4 Achiam et al. (2023). According to the results, a limited number of tokens with extreme reward values (key tokens) dominate the total rewards. For instruction following, the Top-35.6% tokens occupy the highest 80% rewards for chosen responses, while the lowest 37.4% tokens only occupy 20% rewards for rejected responses. In QA, selecting top-50% key tokens can cover over 80% of effective supervision signals. These observations prove that not all tokens are equally effective in preference alignment, and optimizing on all available tokens can be noisy and inefficient Lin et al. (2024); Chen et al. (2024c). Some works explored only optimizing on selected response fragments, but their selection strategies are complex and expensive,such as iterative Monte-Carlo tree search Xie et al. (2024); Chen et al. (2024b) or annotations from human/capable LLMs Lai et al. (2024); Yoon et al. (2024). The above limitations underscore the importance of selective training and efficient token selection strategies in improving preference optimization algorithms.
@@
Based on these intuitions, we propose Selective Preference Optimization (SePO), a novel selective alignment strategy that centers on efficient key token selection without requiring strong, fine-grained supervision signals. We theoretically show that Direct Preference Optimization (DPO) Rafailov et al. (2024c) inherently learns a reward function that decouples the response-level reward values into token level Rafailov et al. (2024b). Based on this conclusion, we propose the first DPO-based token selection method, which trains an oracle model on a moderate-scale subset of the target data distribution, aiming to parameterize an optimal token-level reward function. This token selection strategy has two key advantages: 1) the oracle modeling process is based on the original response-level preference annotations without requiring any extra supervision signals, making it directly applicable to any existing alignment datasets; (2) the cost for token selection can be easily reduced by controlling the oracle model size and the scale of its training subset, which enables cost-efficient selective alignment. We then utilize the estimated reward function to score all tokens within the large-scale target dataset, where tokens with the highest reward values in the chosen responses and tokens with the lowest reward values in the rejected responses are selected as key tokens for alignment. Finally, we design a reference model-free contrastive objective function to optimize the target policy model on the selected tokens. As SePO enables small oracle models to steer selective alignment for much larger policy models, we further explore it as a new paradigm for weak-to-strong generalization Burns et al. (2023), where the focus is two-fold: 1) we leverage weak oracle models to select tokens from in-distribution data for training strong policy models; 2) we train oracle models on out-of-distribution data, which select key tokens to improve target policy model performance and avoid over-optimization Gao et al. (2023); Rafailov et al. (2024a) on the weak supervision data.
@@
  • We propose Selective Preference Optimization, a simple yet effective selective training strategy for preference alignment. It achieves alignment efficiently by obtaining key tokens with an estimated token-level reward function and only optimizing on these selected key tokens;
  • We explore the application of SePO on weak-to-strong generalization. The results show that weak oracle models effectively supervise strong policy models with up to 16.8$\times$ more parameters. SePO also selects useful tokens from weak supervision data to enhance strong policy model performance, alleviating the performance degradation problem commonly encountered in full optimization on out-of-distribution data;
  • We examine SePO on three public evaluation benchmarks. Experimental results show that SePO significantly improves performances on five policy models and outperforms competitive baseline methods by only optimizing 30% key tokens on the target dataset.
Figure 2: The pipeline for Selective Preference Optimization (SePO). SePO mainly consists of three steps: 1) Parameterize a token-level reward function by training an oracle model on a moderate-scale dataset; 2) Score all tokens within the target preference dataset and select key tokens; 3) Training the policy model on selected tokens in a contrastive optimization manner.
@@ 2.1 Preference Alignment of Language Models
$P_{\mathcal{D}}(y_{1}\succ y_{2}|q)=\sigma(r(q,y_{1})-r(q,y_{2}))$ (1)
where $\sigma$ denotes the logistic function and $P_{\mathcal{D}}(y_{1}\succ y_{2})$ denotes the probability that $y_{1}$ is preferred against $y_{2}$. The alignment of language models is typically cast as a KL-constrained optimization problem on the reward values, formalized as
$s.t.\ \mathbb{E}_{q\sim\mathcal{D}}\mathbb{D}_{KL}\left[\pi(y|q)\|\pi_{ref}(y|q)\right]\leq\sigma$
where $\pi$ denotes the aligned policy model, $\pi_{ref}$ denotes the reference policy model. Taking the following Lagrangian, the problem is transformed as:
@@
2.2 Direct Preference Optimization
Ziebart et al. (2008) have shown that Eqn 3 has a closed-form optimal solution, which enables the reward function to be re-parameterized by the optimal policy:
@@ 2.2 Direct Preference Optimization
where $\pi^{*}$ denotes the optimal policy, and $Z(q)$ is the partition function. DPO Rafailov et al. (2024c) bypasses the reward modeling stage by directly substituting this closed-form solution into Eqn. 1, which cancels out $Z(q)$ as it un-changes with the same $q$, yielding the following DPO objective:
@@ 3 Methodology
In this section, we first show DPO training as inherently fitting a token-level reward function that decouples the response-level reward function (Sec. 3.1). Based on this conclusion, we propose Selective Preference Optimization (SePO), which optimizes the target policy model only with selected key tokens by the estimated optimal token-level reward function (Sec. 3.2). We also propose the application of SePO in weak-to-strong generalization (Sec. 3.3). The training pipeline of SePO is shown in Figure 2.
3.1 DPO as Token-level Reward Function Estimator
Due to the auto-regressive nature, the decoding process of LLMs can be naturally formulated as a token-level MDP. The MDP is defined as a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},f,r(s_{t},a_{t}),\rho)$, where $\mathcal{S}$ and $\mathcal{A}$ denote the state space and action space. $s_{t}\in\mathcal{S}$ deontes the current state, consisting of all prompt tokens and current generated tokens (i.e. $s_{t}=\{q|y_{0},...,y_{t}\}$, |||| denotes concatenation). $s_{T}$ denotes the terminal state. $a_{t}\in\mathcal{A}$ denotes the current action, where $\mathcal{A}$ is the token vocabulary. $f$ is the deterministic state transition function that concatenates the current state $s_{t}$ and action $a_{t}$. $\rho$ is an initial state distribution over prompts $q$, where the initial state $s_{1}$ consists of the tokens from $q$. $r(s_{t},a_{t})$ denotes the token-level reward function.
In this section, we show that DPO inherently learns the best estimate on a token-level distribution of the response-level reward values along the token-level MDP trajectory. We begin with the following widely used mild assumption Zhong et al. (2024); Zeng et al. (2024):
@@ 3.1 DPO as Token-level Reward Function Estimator
Any response-level reward function $r$ can be decoupled into the token level as a linear combination of reward values modeled by another token-level reward function $\hat{r}$ along the trajectory for its MDP.
@@
where $s_{t}$, $a_{t}$ are states and actions within the token-level MDP trajectory $\tau=\{s_{1},a_{1},s_{2},...,s_{T}\}$.
With the above assumption, the Bradley-Terry model can be replaced into token level:
$P_{\mathcal{D}}(\tau^{w}\succ\tau^{l}|q)=\sigma(\sum_{i=1}^{N}\hat{r}(s_{i}^{w},a_{i}^{w})-\sum_{j=1}^{M}\hat{r}(s_{j}^{l},a_{j}^{l}))$ (7)
where $\tau^{w}$ and $\tau^{l}$ are the chosen and rejected trajectories (responses), which are assumed to start and end at the same state $s_{1}$ and $s_{T}$.
@@
where $\pi^{*}$ denotes the oracle policy obtained by DPO.
@@
Proof.
This proof is inspired by Rafailov et al. (2024b). Common policy gradient-based RL practices Schulman et al. (2017) optimize Eqn. 3 in token-level MDP with an entropy-bonus $\mathcal{H}(\pi)$ and a KL-constraint with $\pi_{ref}$:
where $V^{*}(s_{T})=0$. Combining Eqn. 10 and 11, we have:
Under Assumption 1, we substitute Eqn. 12 into Eqn. 6, the response-level reward is factorized as follows:
$r(q,\tau)$ $=\sum_{t=1}^{T}r(s_{t},a_{t})$ (13)
Note that in DPO, $V^{*}(s_{1})$ remains unchanged for each response pair as they have the same start state $s_{1}$. This means the preference modeling process for each response pair only depends on the first term of Eqn. 13. Therefore, we conclude that the optimal policy $\pi^{*}$ learned by DPO inherently fits the response-level reward value with another token-level reward function $\hat{r}(s_{t},a_{t})$, which is parameterized as
This indicates our results in Eqn. 8 and completes the proof. ∎
@@ 3.2 Selective Preference Optimization
SePO is guided by the simple idea that "not all tokens are equally effective", which has been widely evaluated Lin et al. (2024); Chen et al. (2024b); Lai et al. (2024). We explore fully utilizing DPO to efficiently select the most useful tokens in modeling human preferences in LLM alignment. Firstly, we train an oracle model on a moderate-scale preference dataset with DPO, aiming to model a token-level reward function for the target data distribution. The reference model is then applied to large-scale data to score all the tokens. The policy model is only trained on selected tokens with highest scores in chosen responses and lowest scores in rejected responses, which are expected as key tokens in achieving alignment.
Oracle Modeling with DPO.
We explore fitting an oracle model with DPO utilizing a moderate-scale dataset via random sampling over the target preference dataset $\mathcal{D}$, which we expect to parameterize a pessimistic estimation of the target reward function.
@@
Proof.
As the reward functions are parameterized via fitting the DPO algorithm on the datasets, we substitute Eqn. 4 into Eqn. 15. As the term $\beta\log\pi_{ref}(y|q)$ and $\beta\log Z(q)$ are unrelated to $\mathcal{S}$, they are easily canceled and we transfer the proof target into comparing the optimal policy functions:
$P_{\mathcal{S}}(X)=\frac{1}{m}\sum_{j=1}^{m}\delta(X=x_{i_{j}})$ (17)
As each $x_{i_{j}}$ is equally likely to be any $x_{i}\in\mathcal{D}$, we have
$\mathbb{E}_{\mathcal{S}}\left[\delta(X=x_{i_{j}})\right]=\frac{1}{N}\sum_{i=1}^{N}\delta(X=x_{i})=P_{\mathcal{D}}(X)$ (19)
Substituting Eqn. 19 into Eqn. 18, we have
Based on the same reference model and empirical data distribution (Eqn. 20), we expect training on $S$ with DPO to obtain an unbiased estimation of the target optimal policy function:
Substituting Eqn. 21 into Eqn. 22, we prove Eqn. 16, which completes the proof. Note that when $m=N$, we have $\mathcal{S}=\mathcal{D}$, and the training process gives an unbiased estimation of target reward function $r_{\mathcal{D}}$. ∎
With the results from Theorem 2, we first perform SFT on the chosen responses of the moderate-scale dataset $\mathcal{S}$ to obtain the base model $\pi_{ref}$:
@@
With the base model, we further perform DPO on the small dataset $\mathcal{S}$ with the objective function Eqn. 5 to obtain the oracle model $\pi_{ora}$. With Theorem 1, we can utilize $\pi_{ref}$ and the oracle model $\pi_{ora}$ to parameterize the optimal token-level reward function for human preferences, which are used for token selections.
@@
We score all tokens within the target preference dataset with the estimated token-level reward function. Based on the token-level MDP and Theorem 1, for each prompt-response pairs $(q,y)$, the score for token $y_{i}$ is calculated as follows:
@@
We define a token selection ratio $k$, which determines the selected proportion for each response. For chosen responses, we utilize the following indicator function for selection:
$\mathbf{I}_{k}^{w}(s(y_{i}))=\begin{cases}1,&\text{if $s(y_{i})$ ranks in highest $k\%$ in $y$}\\ 0,&\text{otherwise}\end{cases}$ (25)
For rejected responses, we change the condition for indicating $1$ to "if $s(y_{i})$ ranks in lowest $k\%$ in $y$" and mark the indicator function as $\mathbf{I}_{k}^{l}(s(y_{i}))$. The intuition behind this action is that key tokens for chosen responses are likely to contribute high token-level rewards, while key tokens for rejected responses are likely with low token-level rewards, whose probabilities are significantly suppressed in $\pi_{ora}$ compared to the base model.
@@
We design a simple contrastive preference optimization objective on the target policy model $\pi_{t}$ with the selected tokens. Specifically, the objective function is designed as follows:
$\mathcal{L}_{SePO}=-\mathbb{E}_{\left(q,y_{w},y_{l}\right)\sim\mathcal{D}}\log\sigma\left(\hat{u}_{w}(q,y_{w})-\hat{u}_{l}(q,y_{l})\right)$ (26)
$\hat{u}_{w}(q,y)=\frac{\gamma}{|y|\cdot k\%}\sum_{i=1}^{|y|}\mathbf{I}_{k}^{w}(s(y_{i}))\log\pi_{t}(y_{i}|q,y_{<i})$
$\hat{u}_{l}(q,y)=\frac{\gamma}{|y|\cdot k\%}\sum_{i=1}^{|y|}\mathbf{I}_{k}^{l}(s(y_{i}))\log\pi_{t}(y_{i}|q,y_{<i})$
where $\gamma$ controls the scaling of the rewards. The SePO objective enables direct adjustment of only crucial tokens for alignment, which avoids noise in both chosen and rejected responses. This design is expected to improve the efficiency and stability of the preference optimization process. The objective is also length-normalized and reference model-free, which prevents bias towards over-length responses and enables memory-efficient alignment Meng et al. (2024); Yuan et al. (2023).
3.3 Application: Weak-to-Strong Generalization
A unique advantage of SePO is that the cost of its token selection process can be easily reduced by controlling the oracle model size, which enables small oracle models to steer the alignment for policy models with much stronger capabilities. Therefore, we explore the application of SePO in weak-to-strong generalization Burns et al. (2023), which aims to elicit the capabilities of strong student models with weak supervision signals. As illustrated in Figure 3, our focus is two-fold:

### v2

@@ Abstract
[…]
Selective Preference Optimization via Token-Level Reward Function Estimation
Kailai Yang1* Zhiwei Liu1  Qianqian Xie2,3 Jimin Huang4 Erxue Min5Sophia Ananiadou1,6 1 National Centre for Text Mining, The University of Manchester 2 School of Artificial Intelligence, Wuhan University 3 Center for Language and Information Research, Wuhan University 4 The Fin AI  5 Baidu Inc. 6 Archimedes Research {kailai.yang,zhiwei.liu,sophia.ananiadou}@manchester.ac.uk {xqq.sincere,erxue.min}@gmail.com;jimin@chancefocus.com
11footnotetext: Work done in collaboration with Baidu Search Science.22footnotetext: Corresponding author.
@@ 1 Introduction
The recent development of large language models (LLMs) has focused on aligning model outputs with human preferences (Ji et al., 2023). During alignment, LLMs are directly optimized on pairwise data and response-level supervision, where popular methods such as reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022; Stiennon et al., 2020) and direct alignment algorithms (Rafailov et al., 2024c; Yuan et al., 2023; Meng et al., 2024) only introduce supervision signals at the end of each response. As deriving preference optimization as bandit problems can lead to sub-optimal solutions and unstable training processes (Zhong et al., 2024; Zeng et al., 2024), many works propose to model LLM decoding as token-level Markov Decision Processes (MDP) and introduce step-wise supervision signals that quantify the value of each action, successfully applied in tasks such as instruction following (Zhong et al., 2024; Yoon et al., 2024) and mathematical reasoning (Xie et al., 2024; Chen et al., 2024b; Lai et al., 2024).
@@
Figure 1: Token-level reward accumulations. As tokens with high rewards are considered key tokens for chosen responses, their Top-K% tokens are accumulated in descending order with the highest rewards. In contrast, rewards are accumulated in ascending orders for rejected responses. More details in Appendix D.
@@
Though achieving outstanding performance, most of these methods are optimized on all available tokens from the training dataset. To validate the effectiveness of this setup, in Figure 1, we present the token-level reward accumulations for 1,000 samples from an instruction following dataset (Cui et al., 2023) where the token-level rewards are assigned by GPT-4 (Achiam et al., 2023). According to the results, a limited number of tokens with extreme reward values (key tokens) dominate the total rewards. The Top-35.6% tokens occupy the highest 80% rewards for chosen responses, while the lowest 37.4% tokens only occupy 20% rewards for rejected responses. These observations prove that not all tokens are equally effective in preference alignment, and optimizing on all available tokens can be noisy and inefficient (Lin et al., 2024; Chen et al., 2024c). Some works explored only optimizing on selected response fragments, but their selection strategies are complex and expensive, utilizing Monte-Carlo Tree Search (Xie et al., 2024; Chen et al., 2024b) or annotations from human/capable LLMs (Lai et al., 2024; Yoon et al., 2024). The above limitations underscore the prospect of selective training and more efficient token selection strategies in improving preference optimization algorithms.
@@
Based on these intuitions, we propose Selective Preference Optimization (SePO), a novel selective alignment strategy that centers on efficient key token selection without requiring strong, fine-grained supervision signals. We show that Direct Preference Optimization (DPO) (Rafailov et al., 2024c) inherently learns a reward function that decouples the response-level reward values into token level (Rafailov et al., 2024b). Based on this conclusion, we propose the first DPO-based token selection method, which trains an oracle model on a moderate-scale subset of the target data distribution, aiming to parameterize an optimal token-level reward function. This token selection strategy has two key advantages: 1) Flexibility: the oracle modeling process is based on the original response-level preference annotations without requiring any extra supervision signals, making it directly applicable to any existing alignment datasets; 2) Efficiency: the cost for token selection can be easily reduced by controlling the oracle model size and the scale of its training subset, which enables cost-efficient selective alignment. We then utilize the estimated reward function to score all tokens within the large-scale target dataset, where tokens with the highest reward values in the chosen responses and tokens with the lowest reward values in the rejected responses are selected as key tokens for alignment. Finally, we design a reference model-free contrastive objective function to optimize the target policy model on the selected tokens.
As SePO enables small oracle models to steer selective alignment for much larger policy models, we further explore it as a new paradigm for weak-to-strong generalization (Burns et al., 2023). Instead of leveraging weak models to provide supervision, 1) we leverage weak oracle models to select tokens from in-distribution data for training strong policy models; 2) we propose to train oracle models on out-of-distribution data, which select key tokens to improve target policy model performance and alleviate over-optimization (Gao et al., 2023; Rafailov et al., 2024a) on the weak supervision data.
@@
  • We propose SePO, the first DPO-based selective training strategy for preference alignment, which applies to any alignment datasets with response-level supervision signals and enables cost-efficient token selection with small-scale oracle models and training data.
[…]
  • We examine SePO on three public evaluation benchmarks. Experiments show that SePO significantly improves performances on six policy models and outperforms competitive baseline methods by only optimizing 30% key tokens with up to 60% reduction in GPU training hours.
Figure 2: SePO mainly consists of three steps: 1) Parameterize a token-level reward function by with a ref-oracle model pair; 2) Select key tokens within the target preference dataset; 3) Train the policy model on selected tokens.
@@ 2 Preliminary
$P_{\mathcal{D}}(y_{1}\succ y_{2}|q)=\frac{exp(r(q,y_{1}))}{exp(r(q,y_{1}))+exp(r(q,y_{2}))}$ (1)
where $P_{\mathcal{D}}(y_{1}\succ y_{2})$ denotes the probability that $y_{1}$ is preferred against $y_{2}$. The alignment of language models is typically cast as a KL-constrained optimization problem on the reward values. Taking the following Lagrangian, the problem is transformed as:
@@
where $\pi$ denotes the aligned policy model, $\pi_{ref}$ denotes the reference policy model.
Direct Preference Optimization
Ziebart et al. (2008) have shown that Eqn 2 has a closed-form optimal solution, which enables the reward function to be re-parameterized by the optimal policy:
@@
where $\pi^{*}$ denotes the optimal policy, and $Z(q)$ is the partition function. DPO (Rafailov et al., 2024c) bypasses the reward modeling stage by directly substituting this closed-form solution into Eqn. 1, which cancels out $Z(q)$ as it un-changes with the same $q$, yielding the following DPO objective:
@@ 3 Methodology
We show DPO as inherently learning the best estimate on a token-level distribution of the response-level reward values (Sec. 3.1). Based on this conclusion, we propose SePO, which optimizes the target policy model with selected key tokens (Sec. 3.2). We also explore SePO as a new paradigm for weak-to-strong generalization (Sec. 3.3). The pipeline is shown in Figure 2.
3.1 Token-level Reward Function Estimator
LLM decoding can be naturally formulated as a token-level MDP, a tuple $(\mathcal{S},\mathcal{A},f,r(s_{t},a_{t}),\rho)$, where $\mathcal{S}$ and $\mathcal{A}$ denote the state and action space. $s_{t}\in\mathcal{S}$ deontes the current state, consisting of all prompt tokens and current generated tokens (i.e. $s_{t}=\{q|y_{0},...,y_{t}\}$, $|$ denotes concatenation). $s_{T}$ denotes the terminal state. $a_{t}\in\mathcal{A}$ denotes the current action, where $\mathcal{A}$ is the token vocabulary. $f$ is the deterministic state transition function. $\rho$ is an initial state distribution over prompts $q$. $r(s_{t},a_{t})$ denotes the token-level reward function.
We begin with the following mild assumption that is widely proven appropriate (Rafailov et al., 2024b; Zhong et al., 2024; Zeng et al., 2024):
@@ 3.1 Token-level Reward Function Estimator
In any token-level MDPs, the reward function $r$ can be decoupled as a linear combination of reward values modeled by another token-level reward function $\hat{r}$ along the trajectory:
@@
with $s_{t}$, $a_{t}$ along the token-level MDP trajectory $\tau=\{s_{1},a_{1},s_{2},...,s_{T}\}$.
With the above assumption, the Bradley-Terry model in Eqn. 1 can be replaced into token level, where the chosen and rejected trajectories are assumed to start and end at the same state.
@@
where $\pi^{*}$ denotes the oracle model obtained via DPO on the reference model.
Proof Sketch.
This proof is heavily inspired by Rafailov et al. (2024b). Starting with the KL-regularized RL objective, the optimization process aims to maximize the expected cumulative reward. Under the maximum entropy RL setting, the optimal policy $\pi^{*}$ is related to the Q-function and value function. The Bellman equation incorporates the KL term, relating the optimal Q-function $Q^{*}$ to the token-level reward $r(s_{t},a_{t})$.
By combining these relationships, the token-level reward can be expressed as:
$r(s_{t},a_{t})=\beta\log\frac{\pi^{*}(a_{t}|s_{t})}{\pi_{ref}(a_{t}|s_{t})}+V^{*}(s_{t})-V^{*}(s_{t+1}),$
where $V^{*}(s_{t})$ is the optimal value function. Under Assumption 1, summing the token-level rewards over all time steps yields the response-level reward. The term $V^{*}(s_{1})$ (the initial state’s value) is constant with the same starting state, which does not affect preference comparisons. Therefore, the preference modeling depends only on the sum of the log-ratio terms. This shows that $\pi^{*}$ inherently aligns an optimal token-level reward function:
$\hat{r}(s_{t},a_{t})=\beta\log\frac{\pi^{*}(a_{t}|s_{t})}{\pi_{ref}(a_{t}|s_{t})},$
which indicates Eqn. 6 and completes the proof. See Appendix C.1 for a detailed proof. ∎
@@ 3.2 Selective Preference Optimization
SePO is guided by the simple idea that "not all tokens are equally effective", which has been widely evaluated (Lin et al., 2024; Chen et al., 2024b; Lai et al., 2024). We explore fully utilizing DPO to efficiently select the most useful tokens in modeling human preferences in LLM alignment. Firstly, we train an oracle model on a moderate-scale preference dataset with DPO, aiming to model a token-level reward function for the target data distribution. The reward function is then applied to large-scale data to score all the tokens. The policy model is only trained on selected tokens with highest scores in chosen responses and lowest scores in rejected responses, which are expected as key tokens in achieving alignment.
Oracle Modeling with DPO
We present the following Theorem to prove the viability of random sampling on the target preference dataset to reduce the cost of oracle model training:
@@
Proof Sketch.
As the reward functions are parameterized via fitting the DPO algorithm, we replace Eqn. 3 into Eqn. 7 and reduce this inequality to comparing the expected optimal policy functions:
$\mathbb{E}_{\mathcal{S}}[\log\pi_{\mathcal{S}}^{*}(y|q)]\leq\log\pi_{\mathcal{D}}^{*}(y|q)$
Since $\mathcal{S}$ is a uniform random sample from $\mathcal{D}$, the empirical distribution $P_{\mathcal{S}}$ is an unbiased estimator of the true distribution $P_{\mathcal{D}}$; that is, $\mathbb{E}_{\mathcal{S}}[P_{\mathcal{S}}(X)]=P_{\mathcal{D}}(X)$. Therefore, training on $\mathcal{S}$ yields an unbiased estimate of the optimal policy: $\mathbb{E}_{\mathcal{S}}[\pi_{\mathcal{S}}^{*}(y|q)]=\pi_{\mathcal{D}}^{*}(y|q)$.
Applying Jensen’s inequality for the concave logarithm function, we have:
$\mathbb{E}_{\mathcal{S}}[\log\pi_{\mathcal{S}}^{*}(y|q)]\leq\log\mathbb{E}_{\mathcal{S}}[\pi_{\mathcal{S}}^{*}(y|q)]=\log\pi_{\mathcal{D}}^{*}(y|q)$
showing that the expected log-optimal policy from $\mathcal{S}$ is less than or equal to that from $\mathcal{D}$ and completes the proof.
See Appendix C.2 for a detailed proof. ∎
Theorem 2 shows that training on a random-sampled subset of the target dataset with DPO can pessimistically estimate the target token-level reward function, and the estimation becomes increasingly accurate with increased sample sizes. With Theorem 2, we first perform SFT on a base model and the chosen responses of the moderate-scale dataset $\mathcal{S}$ to obtain the reference model $\pi_{ref}$:
@@
With the reference model, we further perform DPO on $\mathcal{S}$ with the objective function Eqn. 4 to obtain the oracle model $\pi_{ora}$. With Theorem 1, we can utilize $\pi_{ref}$ and the oracle model $\pi_{ora}$ to parameterize the optimal token-level reward function for human preferences, used for token selection.
@@
We score all tokens within the target preference dataset with the estimated token-level reward function. Based on the token-level MDP and Theorem 1, for each prompt-response pair $(q,y)$, the score for token $y_{i}$ is calculated as follows:
@@
For chosen responses, we utilize the following indicator function for selection:
$\mathbb{I}_{k}^{w}(y_{i})=\begin{cases}1,&\text{if $s(y_{i})$ ranks in highest $k\%$ in $y$}\\ 0,&\text{otherwise}\end{cases}$ (10)
For rejected responses, we change the condition for indicating $1$ to "if $s(y_{i})$ ranks in lowest $k\%$ in yy" and mark the indicator function as $\mathbb{I}_{k}^{l}(y_{i})$. The intuition behind this action (Rafailov et al., 2024b; Zhong et al., 2024) is that key tokens for chosen responses are likely to contribute high token-level rewards, while key tokens for rejected responses are likely with low token-level rewards, whose probabilities are significantly suppressed in $\pi_{ora}$ compared to the reference model.
@@
We design a simple contrastive preference optimization objective (Meng et al., 2024) on the target policy model $\pi_{t}$ with the selected tokens. Specifically, the objective function $\mathcal{L}_{SePO}$ is designed as follows:
$\mathbb{E}_{\left(q,y_{w},y_{l}\right)\sim\mathcal{D}}\log\sigma\left(\hat{u}(q,y_{w},\mathbb{I}^{w}_{k_{w}})-\hat{u}(q,y_{l},\mathbb{I}^{l}_{k_{l}})-\lambda\right)$ (11)
where
$\hat{u}(q,y,\mathbb{I}_{k})=\frac{\gamma}{|y|\cdot k\%}\sum_{i=1}^{|y|}\mathbb{I}_{k}(y_{i})\log\pi_{\theta}(y_{i}|q,y_{<i})$
$\gamma$, $\lambda$ are hyper-parameters, $k_{w}$, $k_{l}$ denote the token selection ratios for chosen/rejected responses, and $\hat{u}$ calculates the length-normalized log-likelihoods for selected tokens. This objective enables direct alignment of only crucial tokens, which avoids noise and prevents bias towards over-length responses (Meng et al., 2024; Yuan et al., 2023).
3.3 SePO for Weak-to-Strong Generalization
A unique advantage of SePO is that its cost can be easily reduced by controlling the base model size, using small oracle models to steer much stronger policy models. We further explore SePO as a new paradigm for weak-to-strong generalization (Burns et al., 2023), which aims to elicit strong student models with weak supervision signals. We propose the following two novel methods (More details in Appendix F):